### v2

# Parametric Traversal for Multi-Dimensional Cost-Aware Graph Reasoning A Policy-Driven Approach for Telecom and Datacenter Infrastructure Graphs

Abstract
Classical path search assumes complete graphs and scalar optimization metrics, yet real infrastructure networks are incomplete and require multi-dimensional evaluation. We introduce the concept of traversal: a generalization of paths that combines existing edges with gap transitions—missing but acceptable connections representing links that can be built. This abstraction captures how engineers actually reason about infrastructure: not just what exists, but what can be realized.
We present a parametric framework that treats planned connections as first-class transitions, scales to large graphs through efficient candidate filtering, and uses multi-dimensional criteria to decide whether a traversal should continue to be explored or be abandoned. We evaluate the framework through representative scenarios in datacenter circuit design and optical route construction in telecommunication networks, demonstrating conditional feasibility, non-scalarizable trade-offs, and policy calibration capabilities beyond the reach of classical formulations.
1 Introduction
Network infrastructures typically consist of physical links between resources (cables connecting devices and ports) and logical links between systems (circuits connecting network nodes or hosts). A path is an ordered sequence of nodes linked together to form an end-to-end connection between a source and a destination. Finding optimal paths to connect devices and nodes is often complex and traditionally relies on error-prone, manual, or cumbersome network tools. Automated path finding has become a key operational capability for digital service providers.
Path search is a foundational problem in graph theory and underlies numerous applications in networking, logistics, robotics, and infrastructure management. Classical formulations typically assume a fully specified graph and aim at optimizing a simple metric such as path length or cumulative weight. However, real-world infrastructure networks rarely conform to these assumptions.
In datacenters and telecommunication networks—whether wired or wireless—infrastructures are often incomplete, heterogeneous, and subject to operational constraints. Engineering decisions are fundamentally about what can be built, not just what exists. Certain connections may be missing but inexpensive to create—a cable within the same rack, a planned fiber extension, a microwave link between two towers—while others are costly or forbidden except in exceptional circumstances. Design tools that ignore this distinction and only traverse existing edges are of limited practical use, as they cannot reason about how to meet new demand when existing connectivity is insufficient.
Furthermore, the notion of an optimal path is highly contextual. Real engineers evaluate feasibility by tracking multiple dimensions simultaneously: the number of new connections to deploy, racks or sites crossed, optical attenuation budgets, cost thresholds. Reducing all these factors to a single scalar weight loses critical information and prevents meaningful optimization.
This paper introduces a programmable path search framework designed to address these challenges. The framework is built around four key design decisions:
Gap transitions as first-class citizens.
The framework models path search as a sequence of transitions between nodes, explicitly distinguishing between edge transitions (existing connections) and gap transitions (missing but acceptable connections). We call such a sequence a traversal—a generalization of the classical notion of path that may combine existing edges with segments yet to be built. This abstraction captures the reality of infrastructure planning where “missing” links are often cheap, planned, or policy-allowed. Traversal should be understood as a planning abstraction rather than a routing artifact.
Separation of acceptability domain and predicate.
To determine which gaps are acceptable, one could evaluate all $O(|N|^{2})$ node pairs—computationally prohibitive for large graphs. Instead, the framework separates the problem into a fast, indexable acceptability domain that restricts candidates (using spatial indices, rack/site scoping, or vendor compatibility rules) and a fine-grained acceptability predicate for detailed evaluation. This mirrors how production systems actually work and enables scaling to large infrastructures.
Multi-dimensional accumulation state.
Rather than scalarizing path costs into a single weight, the framework associates each partial traversal with an arbitrary accumulation state—tracking gap counts, racks crossed, attenuation budgets, or any combination thereof. This preserves the information engineers need to make meaningful decisions.
Demonstrative evaluation on industrial scenarios.
Rather than benchmarking against classical algorithms—which would require reducing the model to a scalar formulation it is designed to transcend—we evaluate the framework through representative scenarios drawn from datacenter and telecommunication network design. These scenarios demonstrate conditional feasibility under business rules, non-scalarizable trade-offs, and ex-ante policy calibration: properties that classical formulations cannot express.
Contributions.
This paper makes the following contributions:
Traversal-level connectivity model.
We introduce a formal model of connectivity based on traversals rather than paths, enabling reasoning over both existing connections and potential deployable transitions in incomplete infrastructure graphs.
Explicit separation of reasoning concerns.
The framework cleanly separates candidate generation (acceptability domain), feasibility filtering (acceptability predicate), resource accumulation, and exploration policies, exposing independent control levers for operational reasoning.
Policy-driven feasibility reasoning.
We show how operational policies (e.g., deployment limits, physical budgets, service tiers) can be encoded as exploration predicates, making feasibility an explicit and inspectable outcome of the traversal process.
Support for non-scalarizable trade-offs.
The traversal model naturally represents multiple admissible solutions that are not totally ordered by a single cost metric, reflecting realistic engineering trade-offs that cannot be captured by classical shortest-path formulations.
Demonstrative evaluation on telco and datacenter scenarios.
Through structurally realistic scenarios, we demonstrate conditional feasibility, non-comparable design alternatives, and policy calibration capabilities that are not expressible with traditional graph-based routing approaches.
2 Related Work
Reasoning about connectivity has long been studied in graph theory and, more recently, in networking research. However, these traditions have evolved largely independently. On the one hand, theoretical work has progressively enriched shortest-path search with multi-objective and resource constraints, while separately studying how to augment graph connectivity. On the other hand, networking research has advanced programmability and intent-based abstractions, but often leaves the feasibility reasoning layer implicit. This section reviews both lines of work and clarifies how the proposed traversal framework relates to them.
2.1 Graph-Theoretic Foundations: From Optimal Paths to Connectivity Augmentation
Shortest paths under constraints.
Classical shortest-path algorithms assume a fully specified graph and optimize a scalar objective. Over time, this formulation has been extended to incorporate increasingly rich constraint structures.
Time-dependent shortest path problems already demonstrate that feasibility and optimality may depend on external policies rather than solely on topology. Dean dean2004timedependent studies minimum-cost paths in time-dependent networks with waiting policies, showing how constraints on waiting behavior alter the structure of feasible paths. In these formulations, however, the graph remains fully defined; policies influence traversal cost or timing, not the existence of transitions.
A major line of research generalizes shortest paths to multi-resource and multi-objective settings. Resource-Constrained Shortest Path (RCSP) problems, surveyed by Pugliese and Guerriero pugliese2013rcsp , associate multiple resource accumulations with each path and rely on dominance-based label-setting or label-correcting methods to prune infeasible or dominated solutions. Multi-Objective Shortest Path (MOSP) problems further extend this perspective by treating costs as vectors and computing Pareto-efficient path sets. Foundational algorithmic developments include dominance-based extensions of Martins’ algorithm gandibleux2006martins , scalable parallel label-setting approaches sanders2013parallel , and more recent heuristic-search variants such as multi-objective D* Lite ren2022modlite . A recent survey by Salzman et al. salzman2023mosp synthesizes progress in heuristic multi-objective search and highlights scalability challenges in high-dimensional cost spaces.
These contributions share two core assumptions: (1) the graph topology is explicitly given; (2) search operates over existing edges only. The main research questions concern optimality, dominance pruning, and scalability under vector-valued costs. In contrast, the traversal framework proposed in this paper retains multi-dimensional accumulation and pruning mechanisms but alters the problem setting itself: transitions are not restricted to existing edges. Instead, they may include acceptable but unrealized connections, whose admissibility is governed by domain-specific policies. The focus shifts from optimal path selection in a fixed topology to conditional feasibility in an incomplete one.
Graph connectivity augmentation.
A distinct branch of graph theory addresses incomplete connectivity through augmentation. Graph augmentation problems study how to add a minimum set of edges to satisfy connectivity properties such as $k$-edge or $k$-vertex connectivity. Early approximation algorithms were developed by Frederickson and Ja’Ja’ frederickson1981augmentation , and comprehensive treatments appear in surveys such as Frank and Jordán frank2015augmentation . In this setting, the objective is global: modify the graph so that it satisfies a structural property with minimal cost.
While graph augmentation explicitly acknowledges missing edges, it treats augmentation as a design decision applied to the entire graph. The output is a modified topology satisfying a target property. In contrast, the traversal framework does not compute a globally augmented graph. Instead, it explores potential transitions locally and conditionally during search. Missing-but-acceptable connections are considered as first-class transitions, yet no permanent modification of the graph is performed. Feasibility emerges at the level of individual traversals under operational constraints, rather than as a global connectivity guarantee.
Thus, the proposed model can be understood as occupying an intermediate position between classical path search (fixed topology) and graph augmentation (global topology modification): it reasons over conditional transitions during search without transforming the underlying graph structure.
2.2 Networking Research: From Programmability to Intent and Feasibility
Programmable and policy-driven networks.
In parallel with theoretical developments, networking research has progressively abstracted control and management. Software-defined networking (SDN) and related paradigms decouple control logic from forwarding behavior, enabling centralized decision-making and policy enforcement. Building on this foundation, intent-based networking (IBN) aims to express high-level operational objectives independently of their implementation.
RFC 9315 rfc9315 provides a formal clarification of intent concepts, distinguishing intent from policy and outlining the functional components required for intent realization. Recent surveys, such as Leivadeas and Falkner leivadeas2023ibn , review architectural patterns for intent translation, policy enforcement, and assurance in automated networks. These works significantly advance the abstraction of network management and formalize how desired states can be expressed and validated.
However, while intent-based frameworks define what the network should achieve, they often assume that the underlying connectivity can be computed over a known topology. The reasoning process that determines whether a connectivity objective is physically realizable—especially in partially documented or incomplete infrastructures—remains embedded within system-specific implementations. Feasibility is typically treated as an outcome of routing or provisioning tools rather than as an explicit modeling construct.
Routing with multiple constraints in operational networks.
Routing with multiple constraints has been extensively studied in the context of QoS provisioning and software-defined networking (SDN). Comprehensive surveys, such as Guck et al. guck2018qos , review delay-constrained and multi-constrained routing algorithms designed for centralized SDN control. These approaches compute feasible paths under delay or bandwidth constraints and improve expressiveness beyond scalar shortest-path routing.
However, they operate over explicitly defined network graphs in which all admissible links are assumed to be known and represented in the topology. In practical infrastructure planning scenarios—such as optical route construction or datacenter cabling—the topology is often incomplete, evolving, or only partially documented. Engineers frequently reason not only about existing links but also about connections that could be deployed under operational constraints. Traditional constraint-based routing formulations do not model this distinction explicitly, as feasibility is evaluated over a fixed topology. In contrast, our framework enables reasoning over conditionally admissible transitions, making connectivity itself policy-dependent rather than purely topological.
2.3 Positioning of the Traversal Framework
The traversal framework proposed in this paper integrates and extends these lines of work along three dimensions.
First, like RCSP and MOSP approaches, it maintains multi-dimensional accumulation states and relies on pruning predicates to control combinatorial growth. However, it does not seek a single optimal solution or even a Pareto frontier over a fixed topology; instead, it enumerates admissible traversals whose feasibility depends on policy-defined constraints.
Second, unlike graph augmentation, it does not compute a globally modified graph satisfying a structural property. Acceptable but unrealized connections are treated as conditional transitions explored during search, without altering the base graph. Connectivity is therefore a property of traversals under policy, not a static property of the graph.
Third, while intent-based networking and policy-driven architectures define high-level objectives and governance mechanisms, the traversal model provides an explicit reasoning layer that bridges intent and physical realizability in incomplete infrastructures. The separation between acceptability domain and predicate, together with externalized accumulation and exploration predicates, makes feasibility an inspectable and programmable construct rather than an implicit by-product of routing.
In summary, prior work has enriched path search with multi-dimensional costs, formalized connectivity augmentation, and abstracted network intent. The traversal framework unifies these perspectives by introducing policy-governed transition generation in incomplete typed graphs, enabling conditional connectivity reasoning that is neither reducible to classical shortest-path optimization nor to global graph augmentation.
3 Parametric Traversal Model
3.1 Typed Graph
Let $\mathcal{P}_{N}$ and $\mathcal{P}_{E}$ be arbitrary sets representing node properties and edge properties, respectively.
Definition 1 (Typed Graph).
A typed graph is a tuple $G=(N,E,\rho_{N},\rho_{E})$ where:
  • NN is a finite set of nodes,
  • $E\subseteq N\times N$ is a set of directed edges,
  • $\rho_{N}:N\rightarrow\mathcal{P}_{N}$ assigns properties to nodes,
  • $\rho_{E}:E\rightarrow\mathcal{P}_{E}$ assigns properties to edges.
The property sets $\mathcal{P}_{N}$ and $\mathcal{P}_{E}$ are application-specific and may encode physical, geographical, or operational characteristics.
3.2 Transitions and Traversals
A central feature of this framework is the ability to traverse not only existing edges but also acceptable connections between nodes that are not yet realized in the graph. We call such acceptable connections gaps.
To determine which connections are acceptable, one could in principle evaluate all $O(|N|^{2})$ pairs of nodes. However, this is computationally prohibitive for large graphs. Moreover, in practice, only a small subset of pairs is relevant.
To address this, we separate the problem into two stages:
  1. 1.
    The acceptability domain restricts, for each node $n$, the set of candidate nodes to consider. This restriction can leverage efficient data structures such as spatial indices, type-based filtering, or precomputed neighborhood structures.
  2. 2.
    The acceptability predicate performs fine-grained evaluation on the restricted set of candidates, encoding complex business rules or constraints.
Definition 2 (Acceptability Domain).
For each node $n\in N$, the acceptability domain $\gamma_{n}\subseteq N$ is the set of nodes to evaluate for acceptability from $n$.
Definition 3 (Acceptability Predicate).
An acceptability predicate is a function
$\alpha:\{(n,m)\in N\times N\mid m\in\gamma_{n}\}\rightarrow\{\mathrm{true},\mathrm{false}\}$
that determines whether a connection between two nodes is acceptable.
Definition 4 (Transition).
Given a typed graph $G=(N,E,\rho_{N},\rho_{E})$, acceptability domains $(\gamma_{n})_{n\in N}$, and an acceptability predicate $\alpha$, a transition is a pair $t=(n,m)\in N\times N$ that is either:
  • an edge transition if $(n,m)\in E$, or
  • a gap transition if $(n,m)\notin E$, $m\in\gamma_{n}$, and $\alpha(n,m)=\mathrm{true}$.
We denote by $T$ the set of all transitions.
Definition 5 (Traversal).
A traversal in a typed graph $G$ is a finite sequence of transitions $\tau_{k}=(t_{1},\ldots,t_{k})$ where $t_{i}=(n_{i},m_{i})$ and $m_{i}=n_{i+1}$ for all $i\in\{1,\ldots,k-1\}$. The integer $k\geq 0$ is the length of the traversal. We denote by $\mathcal{T}$ the set of all traversals.
A traversal is said to be edge-connected if all its transitions are edge transitions; in this case it corresponds to a classical path in the graph. A traversal containing at least one gap transition is called a gapped traversal.
3.3 Accumulation
In classical graph algorithms, paths are often evaluated using a single scalar metric such as length or total weight. However, real-world applications frequently require tracking multiple properties simultaneously: costs, risks, resource consumption, constraint violations, or any combination thereof. Here, costs and risks are understood in a broad sense—they may represent monetary expenses, latency, energy consumption, failure probability, or any domain-specific metric relevant to the application.
The accumulation mechanism addresses this need by associating an arbitrary state with each traversal. This state is computed incrementally as the traversal extends, allowing the search algorithm to:
  • evaluate and compare traversals based on complex, application-specific criteria,
  • prune branches early when the accumulated state indicates infeasibility,
  • return not only the traversal itself but also its associated properties.
Let $\mathcal{A}$ be an application-specific set of accumulation states (e.g., $\mathbb{R}^{+}\times\mathbb{R}^{+}$ for costs and distances).
Definition 6 (Accumulation).
An accumulation over $\mathcal{A}$ is defined by:
  • an initial state $a_{0}\in\mathcal{A}$, and
  • a step function $g:\mathcal{A}\times\mathcal{T}\rightarrow\mathcal{A}$.
The accumulation of a traversal $\tau_{k}$ of length $k$ is denoted $a_{\tau_{k}}$ and defined recursively as:
$a_{\tau_{k}}=\begin{cases}a_{0}&\text{if }k=0\\ g(a_{\tau_{k-1}},\tau_{k})&\text{if }k\geq 1\end{cases}$
The step function receives the previous accumulation $a_{\tau_{k-1}}$ and the traversal up to step $k$. This allows implementations to use only the last transition $t_{k}$ for simple cases, or to access the full history when needed.
3.4 Exploration Predicate
As the traversal extends, the algorithm must decide how to proceed: should it continue exploring, terminate with a valid solution, or prune the current branch? This decision depends on the traversal and its accumulated state, and is application-specific.
Definition 7 (Exploration Predicate).
An exploration predicate is a function
$\sigma:\mathcal{T}\times\mathcal{A}\rightarrow\{\mathtt{continue},\mathtt{terminate},\mathtt{prune}\}$
that determines the action to take for a traversal $\tau$ with accumulation $a_{\tau}$:
  • $\mathtt{continue}$: the traversal is extended by exploring successors,
  • $\mathtt{terminate}$: the traversal is added to the solution set and exploration continues from its successors,
  • $\mathtt{prune}$: the traversal is discarded and its successors are not explored.
4 Traversal Search Algorithm
This section presents a generic algorithm for finding traversals in a typed graph, combining the components defined in Section 3: the typed graph $G=(N,E,\rho_{N},\rho_{E})$, the acceptability domain $\gamma_{n}$, the acceptability predicate $\alpha$, the accumulation function $g$, and the exploration predicate $\sigma$.
The algorithm explores the space of partial traversals. A traversal state is a pair $(\tau,a_{\tau})$ where $\tau\in\mathcal{T}$ is a traversal and $a_{\tau}\in\mathcal{A}$ is its accumulation. For a traversal $\tau_{k}=(t_{1},\ldots,t_{k})$ with $t_{k}=(n_{k},m_{k})$, we call $m_{k}$ the current node.
The algorithm maintains a frontier $\mathcal{F}$: a collection of traversal states awaiting expansion. The frontier policy $\phi$ determines the order in which states are extracted (e.g., FIFO for BFS, priority queue for UCS or A*, bounded-width for beam search).
When a state $(\tau_{k},a_{\tau_{k}})$ with current node $n$ is extracted, successor states are generated:
  • Edge transitions: for each $m$ such that $(n,m)\in E$, form $\tau_{k+1}=\tau_{k}\cdot(n,m)$ with accumulation $a_{\tau_{k+1}}=g(a_{\tau_{k}},\tau_{k+1})$.
  • Gap transitions: for each $m\in\gamma_{n}$ such that $(n,m)\notin E$ and $\alpha(n,m)=\mathrm{true}$, form $\tau_{k+1}=\tau_{k}\cdot(n,m)$ with accumulation $a_{\tau_{k+1}}=g(a_{\tau_{k}},\tau_{k+1})$.
Here $\tau\cdot t$ denotes the concatenation of transition $t$ to traversal $\tau$.
4.1 Algorithm
The complete procedure is given in Algorithm 1.
0: Typed graph $G=(N,E,\rho_{N},\rho_{E})$, start node $n_{0}$
0: Acceptability domains $(\gamma_{n})_{n\in N}$, acceptability predicate $\alpha$
0: Accumulation function $g$, initial accumulation $a_{0}$
0: Exploration predicate $\sigma$, frontier policy $\phi$
0: Set of valid traversals $\mathcal{S}$
[…]
4:while $\mathcal{F}\neq\emptyset$ do
5:  $(\tau,a_{\tau})\leftarrow\textsc{Extract}(\mathcal{F},\phi)$
6:  $s\leftarrow\sigma(\tau,a_{\tau})$
7:  if $s=\mathtt{prune}$ then
8:   continue
9:  end if
10:  if $s=\mathtt{terminate}$ then
11:   $\mathcal{S}\leftarrow\mathcal{S}\cup\{(\tau,a_{\tau})\}$
12:  end if
13:  $n\leftarrow\textsc{CurrentNode}(\tau)$
14:  $V_{\tau}\leftarrow\textsc{VisitedNodes}(\tau)$ {Nodes in $\tau$}
15:  for each $m$ such that $(n,m)\in E$ or ($m\in\gamma_{n}$ and $\alpha(n,m)$) do
16:   if $m\in V_{\tau}$ then
17:    continue {Avoid cycles}
18:   end if
19:   $\tau^{\prime}\leftarrow\tau\cdot(n,m)$
20:   $a_{\tau^{\prime}}\leftarrow g(a_{\tau},\tau^{\prime})$
21:   $\textsc{Insert}(\mathcal{F},(\tau^{\prime},a_{\tau^{\prime}}),\phi)$
22:  end for
23:end while
24:return $\mathcal{S}$
Algorithm 1 Parametric Traversal Search. The algorithm explores traversal states from a frontier, extending partial traversals through edge and gap transitions. The exploration predicate $\sigma$ controls pruning and termination; the accumulation function $g$ tracks multi-dimensional state along each traversal.
The function $\textsc{CurrentNode}(\tau)$ returns the current node: $n_{0}$ if $\tau$ is empty, otherwise the target node of the last transition in $\tau$. The function $\textsc{VisitedNodes}(\tau)$ returns the set of all nodes appearing in the traversal $\tau$.
4.2 Complexity
Let $|N|$ denote the number of nodes and $|E|$ the number of edges. At a node $n$, the algorithm considers at most $\deg(n)+|\gamma_{n}|$ successor transitions, where $\deg(n)$ denotes the out-degree of $n$ in $G$.
In the worst case, the branching factor is bounded by $\max_{n\in N}\deg(n)+\max_{n\in N}|\gamma_{n}|$. Assuming that the algorithm explores all traversals up to a maximum length $L$, and using the average out-degree
$\bar{d}=\frac{1}{|N|}\sum_{n\in N}\deg(n)=\frac{|E|}{|N|}$
as an approximation, the total number of traversal states explored is bounded by
$O\bigg(\Big(\frac{|E|}{|N|}+\max_{n\in N}|\gamma_{n}|\Big)^{L}\bigg),$
which highlights the combined impact of graph density and acceptability domain size on the search complexity. Note that the cycle avoidance mechanism (line 11 of Algorithm 1) ensures that each traversal visits at most $|N|$ distinct nodes, implicitly bounding $L\leq|N|$.
This exponential complexity makes the exploration predicate $\sigma$ and the accumulation function $g$ critical for practical performance. These functions encode domain-specific knowledge about what constitutes a viable traversal: the accumulation tracks relevant metrics (cost, resource consumption, constraint violations), while $\sigma$ determines when to prune, terminate, or continue.
Crucially, $\sigma$ and $g$ are external to the algorithm—they encapsulate business rules and domain expertise. A well-designed accumulation that captures the essential properties of traversal quality, combined with an exploration predicate that aggressively prunes unpromising branches, can reduce the effective complexity by orders of magnitude. Conversely, a poorly specified $\sigma$ that fails to prune early will result in exhaustive exploration and intractable runtime.
The framework thus places the responsibility for efficiency on the domain expert: the algorithm provides the exploration mechanism, but the user must define what makes a traversal good or bad through $g$ and $\sigma$. The framework does not aim at improving worst-case complexity bounds over classical graph search. Its contribution lies in making explicit where combinatorial growth originates and how domain knowledge can be used to control it.
4.3 Termination
Proposition 1 (Termination under bounded accumulation).
Assume that:
  1. 1.
    the graph $G=(N,E)$ is finite,
  2. 2.
    for each node $n\in N$, the acceptability domain $\gamma_{n}$ is finite,
  3. 3.
    there exists a bound $L\in\mathbb{N}$ such that $\sigma(\tau,a_{\tau})=\mathtt{prune}$ for all traversals $\tau$ with length greater than $L$.
Then Algorithm 1 terminates after a finite number of iterations.
Proof.
Since $N$ and all acceptability domains $\gamma_{n}$ are finite, the number of possible successors of any traversal state is finite. Under assumption (3), no traversal of length greater than $L$ is expanded.
Therefore, the algorithm explores only traversals of length at most $L$. The number of such traversals is finite, bounded by
$\sum_{k=0}^{L}\prod_{i=1}^{k}(\deg(n_{i})+|\gamma_{n_{i}}|),$
where $\deg(n_{i})$ denotes the out-degree of node $n_{i}$.
As a result, the frontier becomes empty after finitely many expansions, and the algorithm terminates. ∎
Remark 1.
In practice, the bound on traversal length is typically enforced indirectly through the accumulation function and the exploration predicate, for instance by limiting cost, resource consumption, or the number of gap transitions. This allows termination to be guaranteed without imposing an explicit depth limit.
5 Demonstrative Evaluation
This work is not evaluated through classical performance metrics, as it does not propose a new path-finding algorithm nor claim improved optimality. Instead, we focus on the expressiveness and structural properties of the traversal model, supported by formal analysis and representative case studies.
Our evaluation presents three scenarios that expose the expressive limits of classical formulations and demonstrate how traversal-level reasoning enables conditional feasibility, non-scalarizable trade-offs, and policy calibration. Each scenario carries a distinct message that cannot be reduced to a benchmark comparison.
5.1 Industrial Context
We ground the evaluation in two infrastructure domains.
Datacenter. Rooms contain rows, rows contain racks, and racks contain devices. Client racks house servers and patch panels; distribution racks house switches with upstream connectivity. Structured cabling interconnects patch panels along each row toward distribution racks. Within a rack, cross-connects are inexpensive to deploy; inter-rack cabling is costly and typically avoided. The routing problem is to connect a customer server to upstream network equipment, potentially through gaps representing cables to be deployed or device reconfigurations to be performed (e.g., configuring a VLAN to bridge two ports on a switch).
Telecommunication network. Sites contain optical distribution frames (ODFs), splice boxes, and amplifiers, interconnected by fiber segments. Each fiber segment carries physical properties (length, attenuation). Intra-site extensions are relatively inexpensive; inter-site gaps require civil works. The routing problem is to establish an optical path between two ODFs while respecting physical constraints such as attenuation budgets and distance limits.
Broader applicability. Although the scenarios presented here focus on wired infrastructures, the traversal model applies equally to wireless domains. In microwave radio-link networks, gap transitions can represent planned radio links subject to line-of-sight and frequency licensing constraints. In 5G RAN deployments, they can model potential fronthaul or midhaul connections between distributed units and central units. In satellite networks, gaps may represent inter-satellite or ground-to-satellite links subject to orbital geometry and link budget constraints. In each case, the acceptability domain restricts candidate links by physical feasibility, and the accumulation tracks domain-specific metrics such as signal-to-noise ratio, latency, or spectral efficiency.
5.2 Conditional Feasibility in Incomplete Networks
The first evaluation demonstrates that feasibility is not a binary property of the graph, but a conditional property of a traversal under business rules.
5.2.1 Telco — Optical Route under Physical Constraints
Consider connecting a source ODF to a target ODF in a partially documented network. The following operational rules apply:
  • Intra-site extensions (gaps) are authorized if the distance between endpoints is less than 100 m.
  • Total attenuation must not exceed 30 dB.
  • At most 1 gap transition is allowed.
  • Amplifiers may be traversed at most once.
These rules are not independently scalarizable: attenuation depends on fiber type and length, the gap limit is a hard count constraint, and the amplifier rule introduces a history-dependent condition. No single weighted shortest-path formulation can capture all four simultaneously without losing constraint semantics.
Traversal parametrization.
  • $\gamma_{n}=\{m\in N\mid\mathrm{site}(m)=\mathrm{site}(n)\}$: candidates restricted to the same site.
  • $\alpha(n,m)=\mathrm{true}$ iff $\mathrm{dist}(n,m)<100\,\text{m}$ and fiber types are compatible.
  • $\mathcal{A}=\mathbb{R}^{+}\times\mathbb{R}^{+}\times\mathbb{N}$ with $a_{\tau}=(\text{total length},\text{total attenuation},\text{number of gaps})$.
  • $\sigma(\tau,a_{\tau})=\mathtt{prune}$ if attenuation $30\,\text{dB}$ or gaps >1>1; $\mathtt{terminate}$ if the current node is the target ODF; $\mathtt{continue}$ otherwise.
Result.
A valid traversal is found that includes one intra-site gap, respects the attenuation budget, and satisfies the amplifier constraint. The traversal is returned with its full accumulation state, making the engineering rationale explicit: the gap location, the attenuation contribution of each segment, and the remaining budget are all visible in the output.
This routing problem cannot be expressed as a single weighted shortest path without losing constraint semantics. The traversal framework solves it without post-processing and without constraint relaxation.
5.2.2 Datacenter — Physical Circuit under Client Policy
Consider connecting a customer server to an upstream switch. The following rules apply:
  • Cross-connects (gaps) are authorized within the same rack.
  • Inter-rack gaps are forbidden for standard clients.
  • At most 2 new cables may be deployed.
  • At most 1 row change is allowed.
Traversal parametrization.
  • $\gamma_{n}=\{m\in N\mid\mathrm{rack}(m)=\mathrm{rack}(n)\text{ and }\mathrm{type}(m)\in\{\text{patch panel}\}\}$: candidates restricted to the same rack.
  • $\alpha(n,m)=\mathrm{true}$ iff both $n$ and $m$ have available ports.
  • $\mathcal{A}=\mathbb{N}\times\mathbb{N}\times\mathbb{N}$ with $a_{\tau}=(\text{gaps},\text{racks traversed},\text{rows crossed})$.
  • $\sigma(\tau,a_{\tau})=\mathtt{prune}$ if gaps >2>2 or rows >1>1; $\mathtt{terminate}$ if $\mathrm{upstream}(m_{k})=\mathrm{true}$; $\mathtt{continue}$ otherwise.
Result.
Under a premium client policy (inter-rack gaps allowed, up to 5 gaps), a valid traversal is found. Under a standard client policy (rack-local gaps only, at most 2 gaps), the same query yields no admissible traversal.
The infrastructure is identical in both cases. The feasibility depends entirely on the policy encoded in $\gamma$, $\alpha$, and $\sigma$—not on the graph structure. This distinction is invisible to any formulation that reduces feasibility to graph connectivity.
5.3 Non-Scalarizable Trade-offs in Physical Circuit Design
The second evaluation demonstrates that certain engineering decisions cannot be ordered by a single scalar cost function.
Scenario.
Consider a datacenter where a server-to-upstream connection admits two valid traversals:
Gaps Racks traversed Rows crossed
Traversal $\tau_{1}$ 0 2 1
Traversal $\tau_{2}$ 1 (intra-rack) 1 0
Both traversals satisfy all operational constraints. Neither dominates the other:
  • $\tau_{1}$ requires no new cable but crosses a row boundary;
  • $\tau_{2}$ stays within a single rack and row but requires deploying one cross-connect.
No fixed weighting of these three dimensions can consistently reflect all possible operational policies. An operator prioritizing deployment cost would prefer $\tau_{1}$; an operator prioritizing spatial locality would prefer $\tau_{2}$. Collapsing the accumulation into a scalar erases this distinction.
What the framework provides.
The traversal search returns both $\tau_{1}$ and $\tau_{2}$ as elements of the solution set $\mathcal{S}$, each annotated with its full accumulation state. The final selection can be deferred to a human operator or resolved by a downstream policy function operating on the structured accumulation.
The output of traversal search is a decision space, not a single optimum. This is precisely the regime where the multi-dimensional accumulation model provides value over scalar-cost formulations.
5.4 Policy Calibration and Severity Analysis
The third evaluation demonstrates that the framework can serve as an instrument for calibrating operational policies before deployment.
Scenario.
A telecommunication operator wants to determine the attenuation budget $B$ at which the network becomes predominantly connectable. The question is:
“Starting from what attenuation budget does the majority of ODF pairs become reachable?”
This is not a path-finding query but a structural analysis of the policy itself.
Procedure.
The exploration predicate is parameterized by $B$:
$\sigma_{B}(\tau,a_{\tau})=\mathtt{prune}\quad\text{if total attenuation}>B$
For each value of $B\in\{20,25,30,35\}\,\text{dB}$, we analyze which source-target ODF pairs admit at least one valid traversal. All other parameters ($\gamma$, $\alpha$, $g$) remain fixed.
Result (qualitative).
Budget $B$ Connectivity
$B<25\,\text{dB}$ Very few ODF pairs are reachable
$B\approx 30\,\text{dB}$ Majority of pairs become connectable
$B>35\,\text{dB}$ Marginal gain in additional connectivity
This analysis reveals a phase transition in network connectivity as a function of policy severity. The operator can identify the budget threshold beyond which relaxing the constraint yields diminishing returns, and calibrate the operational policy accordingly—before any physical deployment. This analysis is independent of any specific query workload and characterizes the interaction between topology and policy itself.
The framework thus acts as an instrument for policy analysis, not as an opaque solver returning a single answer.
6 Discussion
The parametric traversal model generalizes classical path search by introducing two key innovations: (1) gap transitions that allow reasoning about missing but acceptable connections, and (2) a fully parametric search where accumulation, pruning, termination, and frontier policy are externalized. This section interprets the main design choices, discusses their trade-offs, and identifies the settings where the framework is—and is not—appropriate.
6.1 Interpretation of the Model
The traversal model shifts the focus from finding optimal paths to reasoning about realizable connectivity. In classical formulations, the graph is a given and the algorithm searches for the best route. Here, the graph is incomplete, and the algorithm explores what could exist under operational constraints. The output is not a single optimum but a structured decision space: a set of traversals, each annotated with its multi-dimensional accumulation state, that captures the trade-offs engineers actually face.
This interpretation positions the framework as a planning instrument rather than a routing engine. It is designed to answer questions such as “can this connection be realized under these constraints?” and “what are the alternatives?”, rather than “what is the shortest path?”.
6.2 Design Choices
Two-stage acceptability.
The separation between the domain $\gamma_{n}$ and the predicate $\alpha$ reflects a distinction between structural invariants and operational state. The domain $\gamma_{n}$ captures stable structural relations—for instance, from a server to patch panels in the same rack—independent of runtime conditions. The predicate $\alpha(n,m)$ evaluates whether a structurally admissible transition is operationally valid at a given time, such as whether a patch panel is available or functional. While $\alpha$ could technically be folded into $\gamma_{n}$, maintaining this separation prevents structural adjacency from being conflated with transient operational constraints, which is essential in infrastructure reasoning.
Frontier policy.
The frontier policy $\phi$ does not merely determine exploration order; it defines the decision semantics of traversal. By parameterizing frontier selection, different notions of preferred solutions emerge from the same structural and policy constraints—prioritizing hop count emphasizes structural proximity, while prioritizing accumulated cost highlights operational efficiency. This separation between transition generation, admissibility, and exploration strategy keeps the traversal engine agnostic to the optimization criterion while preserving the multi-dimensional cost structure. In this sense, $\phi$ acts as a decision layer rather than a fixed algorithmic choice.
Externalized control.
The accumulation function $g$ and the exploration predicate $\sigma$ are deliberately external to the algorithm. They encapsulate business rules and domain expertise, making the framework a generic exploration mechanism that can be specialized without modification.
6.3 Trade-offs
Flexibility vs. complexity.
The framework’s flexibility comes at the cost of increased search complexity. As shown in Section 4.2, the worst-case complexity is $O((|E|/|N|+b)^{L})$, where $b=\max_{n}|\gamma_{n}|$ is the maximum acceptability domain size and $L$ the maximum traversal length. Controlling combinatorial growth depends on the entire parametrization. The domain $\gamma_{n}$ bounds the branching factor; the predicate $\alpha$ further filters candidates within that domain; the accumulation $g$ tracks the state on which $\sigma$ relies to prune early. A weak choice in any of these components—an overly broad domain, a permissive predicate, or an accumulation that fails to capture relevant constraints—can lead to intractable exploration.
Expressiveness vs. optimality.
By preserving multi-dimensional accumulation states, the framework avoids the information loss inherent in scalarization. However, this means it does not produce a single optimal solution in the classical sense. Instead, it returns a Pareto-like set of admissible traversals, deferring the final selection to downstream decision-making.
6.4 When Traversal Is and Is Not Appropriate
The framework is designed for settings where connectivity is uncertain, conditional, or partially realized—where the question is not “what is the best route?” but “what can be built, and under what conditions?”. It adds value precisely when the graph is incomplete and feasibility depends on operational policies rather than on topology alone.
In environments where the topology is fully documented and stable—such as pure IP routing over well-provisioned networks—classical shortest-path algorithms remain simpler, faster, and entirely sufficient. The traversal model is not a replacement for Dijkstra or Bellman-Ford; it addresses a different class of problems that these algorithms were not designed to handle.
7 Limitations
Context-independent acceptability.
The acceptability predicate $\alpha(n,m)$ is context-independent: it does not consider the traversal history. A context-dependent variant $\alpha(\tau,n,m)$ would enable richer constraints (e.g., “this gap is acceptable only if no other gap has been used in the last three transitions”) but at increased computational cost and modeling complexity.
No formal optimality guarantees.
The framework provides termination guarantees under bounded accumulation (Proposition 1) but does not guarantee optimality in the classical sense. Since the accumulation space $\mathcal{A}$ is arbitrary and may not admit a total order, the notion of “optimal traversal” is not always well-defined. The framework is designed to enumerate admissible solutions, not to certify global optimality.
Scalability under weak pruning.
The exponential worst-case complexity means that poorly designed exploration predicates can lead to intractable runtimes. The framework places the responsibility for efficiency on the domain expert: without aggressive pruning through $\sigma$, the search may degenerate into exhaustive enumeration.
Single-traversal reasoning.
The current model reasons about individual traversals independently. It does not natively support constraints that span multiple traversals, such as shared capacity limits or correlated failure domains. This limits its applicability to scenarios requiring joint optimization over sets of traversals.
Evaluation scope.
The demonstrative evaluation is based on representative scenarios rather than large-scale empirical benchmarks. While this is consistent with the paper’s goal of demonstrating expressiveness rather than computational performance, it leaves open the question of practical scalability on production-sized infrastructures.
8 Future Work
From single traversals to redundant structures.
In practice, infrastructure connectivity often requires redundancy: multiple disjoint traversals between endpoints to ensure fault tolerance, maintenance flexibility, or service-level guarantees. The current framework focuses on discovering individual traversals, whereas real deployments may require pairs of disjoint traversals, rings, or more complex topologies. Extending the model to reason about sets of traversals—while enforcing shared constraints across them (e.g., shared risks, common physical segments, or correlated costs)—is a natural direction for future work.
Traversal sets and global constraints.
The current model evaluates feasibility at the level of individual traversals, with costs accumulated independently. In many operational scenarios, however, feasibility depends on global constraints spanning multiple traversals, such as shared capacity limits, common failure domains, or cumulative deployment budgets. An important extension would be to lift cost accumulation and policy evaluation from individual traversals to traversal sets, enabling reasoning over collective feasibility and interactions between concurrent connectivity decisions.
Integration with industrial inventory and intent models.
While this work remains independent of any specific data model, it naturally aligns with industrial standards such as TM Forum ODA and SID, which provide structured representations of network resources and relationships. Future work includes instantiating the traversal framework directly on top of such models and connecting it to intent-based specifications, thereby enabling an explicit reasoning layer between high-level intents and realizable infrastructure actions.
Traversal as a reasoning substrate for agentic systems.
Beyond infrastructure planning, the parametric traversal framework may serve as a structural reasoning component within agentic decision systems. Its explicit separation of structural generation, policy validation, and exploration strategy makes it suitable as a controllable and interpretable planning layer. An agent could leverage traversal search to evaluate connectivity hypotheses, simulate policy adjustments, or explore counterfactual scenarios without collapsing multi-dimensional costs into scalar objectives. Future work may investigate how traversal-based reasoning integrates into higher-level agent architectures, enabling structured decision-making over incomplete graphs while preserving explicit policy control and cost transparency. This perspective positions traversal not merely as a search mechanism, but as a symbolic reasoning primitive over structured environments.
Hybrid reasoning and approximation strategies.
The framework deliberately prioritizes expressiveness over worst-case complexity guarantees. A promising research direction is the integration of heuristic, approximate, or learning-assisted strategies to guide traversal exploration, while preserving the explicit policy and cost structure of the model. This could enable scalable reasoning in larger infrastructures without collapsing the multi-dimensional decision space into a single scalar objective.
Practical use cases across infrastructure domains.
The demonstrative evaluation presented in this paper focuses on wired physical-layer scenarios in datacenter and fiber-optic networks. A natural continuation is a series of practical studies exploring the framework across a broader range of infrastructure domains. On the physical layer, this includes wireless domains such as microwave radio-link planning, 5G RAN fronthaul and midhaul design, and satellite link engineering, where gap transitions model potential radio or optical free-space links subject to propagation, licensing, and link budget constraints. Beyond the physical layer, the framework can also be applied to logical-layer connectivity: data-link circuits, VLAN paths, MPLS label-switched paths, or overlay tunnels, where gaps represent provisionable logical connections and the accumulation tracks bandwidth, latency, or hop-count budgets. Such a series of studies would validate the generality of the traversal model and identify domain-specific parametrizations that emerge in practice.
9 Conclusion
We presented a parametric traversal framework for typed graphs that addresses the fundamental gap between classical path search algorithms and the realities of infrastructure planning. The framework makes four key contributions:
Gap transitions as a practical abstraction.
By treating missing-but-acceptable connections as first-class transitions, the framework captures the essence of infrastructure design: engineering decisions are about what can be built, not merely what exists. This abstraction reflects how datacenters and telecommunication networks actually evolve, where deploying a new cable or extending a fiber path is part of normal operational planning.
Scalable acceptability evaluation.
The separation of acceptability domain and predicate mirrors production-grade system design. Fast, indexable candidate restriction (spatial indices, rack scoping, vendor compatibility) combined with fine-grained predicate evaluation enables the framework to scale to large infrastructures without sacrificing expressiveness.
Preservation of multi-dimensional state.
By maintaining arbitrary accumulation states rather than collapsing metrics into a single scalar, the framework preserves the information needed for meaningful decisions—whether physical metrics, operational constraints, or any domain-specific criteria.
Expressiveness demonstrated through case studies.
The demonstrative evaluation (Section 5) shows that the framework addresses routing problems that resist classical formulations: conditional feasibility under business rules, non-scalarizable trade-offs between competing dimensions, and ex-ante policy calibration. These scenarios are drawn from datacenter and telecommunication network design and reflect actual operational challenges encountered in production environments.